# Wealth Thermalization Hypothesis

### v2

@@ Abstract
We introduce the wealth thermalization hypothesis according to which the wealth shared in a country or the whole world is described by the Rayleigh-Jeans thermal distribution with two conserved quantities of system wealth and norm or number of agents. This distribution depends on a dimensional parameter being the ratio of system total wealth and its dispersion range determined by highest revenues. At relatively small values of this ratio there is a formation of the Rayleigh-Jeans condensate, well studied in such physical systems as multimode optical fibers. This leads to a huge fraction of poor households and a small oligarchic fraction which monopolizes a dominant fraction of total wealth thus generating a strong inequality in human society. We show that this thermalization gives a good description of real data of Lorenz curves of US, UK, the whole world and capitalization of S&P500 companies at New York Stock Exchange. Possible actions for inequality reduction are briefly discussed.
@@
Introduction. The wealth distribution in the human society is characterized by a striking inequality (see e.g. piketty1 ; piketty2 ; boston ). Thus for the whole world 50% of the population owns only 2% of total wealth, while 10% of population owns 75% of total wealth and 1% of population owns 38% of total wealth piketty2 .
@@
The distribution of wealth is usually described by the Lorenz curve lorenz ; boston which gives the dependence of cumulated normalized wealth $0\leq w\leq 1$ on the cumulated normalized fraction of population or households $0\leq h\leq 1$. Thus the equipartition of wealth corresponds to the diagonal $w=h$ and the doubled area between diagonal and the Lorenz curve $w(h)$ determines the Gini coefficient $0\leq G\leq 1$ gini ; boston . Values of $G$ can be found in wikigini for world countries in 2021 being in the range $0.900.59<G<$ 0.90; for the whole world $G=0.889$.
@@
Diverse methods of statistical mechanics and physical kinetics mayer ; landau ; lifshitz have been proposed and used by different research groups angle ; redner ; yakovenko1 ; bouchaud ; yakovenko2 ; chakraborti ; boghosian1 ; boghosian2 ; perotti . Various models of interacting agents are investigated including Random Asset Exchange models angle ; redner ; yakovenko1 ; bouchaud ; yakovenko2 ; chakraborti ; boghosian1 ; boghosian2 ; perotti . In several of these models there is appearance of some kind of oligarchic phase with a significant wealth accumulation by a group of agents bouchaud ; boghosian1 ; boghosian2 ; perotti . The specific arguments are presented in a favor of the Boltzmann-Gibbs type description of distribution of money, wealth and income yakovenko1 ; yakovenko2 . Also a nonlinear Fokker–Planck description of asset exchange is proposed boghosian1 ; boghosian2 with emergence of oligarchic phase. A few important elements are stressed in boghosian1 ; boghosian2 : the conservation of two integrals of system evolution being the total wealth and total norm (or number of agents), the argument in favor of consideration of wealth instead of money based on the small-transaction approximation.
The above models give interesting insights for understanding of certain features of wealth distribution in the world countries but they remain model specific and their universality remains questionable. The universality of the Boltzmann-Gibbs thermal distribution is the ground element of the approach developed in yakovenko1 ; yakovenko2 but it does not capture emergence of a huge condensate of poverty in various countries.
@@
Here we assume that the system wealth has certain states $0\leq m<N$ with energies $E_{m}$ and the population probabilities in these states are $\rho_{m}$. Also $T$ is the system temperature and $\mu(T)$ is the chemical potential dependent on $T$. As in boghosian1 there are two conserved integrals of motion being the total norm of population, fixed to be unity for convenience, $\sum_{m}\rho_{m}=1$, and the system average wealth being its total energy $\sum_{m}E_{m}\rho_{m}=E$. For a given system energy $E$ and unity norm these two integrals of motion determine the system temperature $T$ and its chemical potential $\mu(T)$. The entropy $S$ of the system is determined by the usual relation mayer ; landau : $S=-\sum_{m}\rho_{m}\ln\rho_{m}$ with the implicit theoretical dependencies on temperature $E(T)$, $S(T)$, $\mu(T)$.

### v3

# received: September 21, 2025

@@ Abstract
In 1955 Fermi, Pasta, Ulam and Tsingou performed first numerical studies with the aim to obtain the thermalization in a chain of nonlinear oscillators from dynamical equations of motion. This model happend to have several specific features and the dynamical thermalization was established only later in other studies. In this work we study more generic models based on Random Matrix Theory and social networks with a nonlinear perturbation leading to dynamical thermalization above a certain chaos border. These systems have two integrals of motion being total energy and norm so that the theoretical Rayleigh-Jeans thermal distribution depends on temperature and chemical potential. We introduce the wealth thermalization hypothesis according to which the society wealth is associated with energy in the Rayleigh-Jeans distribution. At relatively small values of total wealth or energy there is a formation of the Rayleigh-Jeans condensate, well studied in physical systems such as multimode optical fibers. This condensation leads to a huge fraction of poor households at low wealth and a small oligarchic fraction which monopolizes a dominant fraction of total wealth thus generating a strong inequality in human society. We show that this thermalization gives a good description of real data of Lorenz curves of US, UK, the whole world and capitalization of companies at Stock Exchange of New York SE (NYSE), London and Hong Kong. It is also shown that above a chaos border the dynamical Rayleigh-Jeans thermalization takes place also in social networks with the Lorenz curves being similar to those of wealth distribution in world countries. Possible actions for inequality reduction are briefly discussed.
Introduction
In 1872 Ludwig Boltzmann published the fundamental work [1] that became the foundation of the theory of statistical physics and thermalization emerging from the dynamical laws of time reversible classical motion of many-body systems. Of course, certain approximations had been used there, thus considering only pair collisions of gas particles. But this was only the first step in the demonstration that the statistical laws follow from the dynamical equations. One of the important results obtained in [1] was the Boltzmann H-theorem that a system entropy is monotonically growing with time or remains constant at the steady-state.
The first numerical experiment with the aim to directly obtain statistical thermalization from dynamical equations of motion was done by Fermi, Pasta, Ulam and Tsingou in 1955 studying the dynamics of a chain of nonlinear oscillators on the most powerful MANIAC I computer (at this time) with an expectation to obtain the thermal energy equipartition between oscillator modes [2]. However, opposite to expectations the conclusion of authors was that “The results show very little, if any, tendency toward equipartition of energy between the degrees of freedom.” [2].
Several explanations had been proposed to explain this result. Zabusky argued in [3] that in the continuum limit the Fermi-Pasta-Ulam (FPU) problem is close to the Korteweg-de Vries equation with stable solitons shown to be completely integrable [4], as well as the nonlinear Schrödinger equation [5]. Also, at weak nonlinearity the FPU $\alpha$-model is close to the completely integrable Toda lattice [6, 7]. Another explanation of absence of thermalization in the FPU problem was given in [8, 9, 10] showing that below a certain chaos border, determined by the strength of the nonlinear interactions between oscillators, the system is located in the regime of Kolmogorov-Arnold-Moser (KAM) integrability and only above this border an overlap of nonlinear resonances takes place with emergence of chaos and thermalization. Indeed, above a chaos border numerical simulations demonstrated an emergence of dynamical thermalization with energy equipartition as reported in [9, 10]. Possibilities of low energy chaos in the FPU model were discussed in [11],
It should be pointed out that while impressive mathematical results and theorems were obtained by mathematicians (see e.g. [12, 13] and Refs. therein) they remained usually not applicable to thermalization in physical nonlinear systems which usually have a divided phase space (see [14, 15]) where integrable islands of stable motion are often embedded in a chaotic component. Thus the KAM theorem is valid for unrealistically weak nonlinear perturbations [12, 13] and it is more appropriate to use the Chirikov criterion of overlapping resonances to estimate more realistic parameters for a chaos border [14, 15] (even if this criterion is not working for completely integrable systems like the Toda lattice for example).
An overview of the full richness of the FPU model and various regimes of its nonlinear dynamics has been presented 50 years after [2] in the book [16]. The variety of studies presented there clearly demonstrates that this model has an important role in the investigations of nonlinear dynamics. However, at the same time the variety of different features of FPU dynamics indicates that the FPU model does not belong to a class of generic oscillator systems with nonlinear interactions.
With the aim to capture the generic features of dynamical thermalization the nonlinear random matrix model (NLIRM) was proposed in the work [17], submitted 150 years after the Boltzmann article [1]. In this model, the linear unperturbed Hamiltonian is described by a random matrix that can be also viewed as a system of linear oscillators with complex linear couplings. The chaos in this system is induced only by a nonlinear perturbation that can be local or can have a certain interaction range. Thus in this model the unperturbed properties of eigenmodes and eigenenergies are described by the generic Random Matrix Theory invented by Wigner for a description of the spectra of complex nuclei, atoms, and molecules in many-body quantum mechanics [18]. Indeed, RMT finds a variety of applications in multiple areas of physics [19, 20] including systems of quantum chaos where the dynamics is chaotic in the classical limit [21, 22].
In [17] it was shown that above a certain chaos border the dynamical thermalization takes place leading to the Rayleigh-Jeans thermal distribution (see Eq. (I.1 below) well known in classical thermodynamics [23, 24]. However, the dynamics of the NLIRM system has two integrals of motion being the total energy and norm (or probability that is very natural for quantum evolution). Due to this the Rayleigh-Jeans distribution is characterized by temperature and chemical potential. Such a situation appears in various classical systems including dynamics of nonlinear waves (see e.g. [25]). In fact this type of thermal distribution was described and experimentally observed for light propagation in multimode optical fibers (see e.g. review [26] and Refs. below). The important feature of the Rayleigh-Jeans thermalization is the condensation of a main fraction of system norm or probability at the lowest energy modes of the system. However, in [26] the emergence of such thermalization and condensation was attributed to the turbulence like energy flows similar to those of the Kolmogorov-Zakharov turbulence spectra of nonlinear waves [25] (even if it is stated [26] that the dynamics is Hamiltonian). In contrast, it is argued in [17, 27] that such Rayleigh-Jeans thermalization and condensation appear due to dynamical chaos emerging above a chaos border while below this border the thermalization is absent and the system is located in the integrable KAM regime. Certain similarities of this condensation with the Fröhlich condensate proposed for molecular systems at room temperature [28, 29] are discussed in [27]. The striking applications of Rayleigh-Jeans thermalization are discussed in this work.
This work is composed of two parts. In the first Part I, we push forward the Wealth Thermalization Hypothesis according to which the wealth is associated with system energy and the wealth distribution in the human society is described by the Rayleigh-Jeans thermalization and condensation which are at the origin of strong inequality in the world. The comparison of the obtained thermalization results with the real data of wealth inequality in countries and stock exchange markets confirms the validity of the Rayleigh-Jeans thermal description. In the second Part II, we provide certain additional arguments and justifications for this hypothesis. In particular, we argue that social networks, actively investigated in the network science and society (see e.g. books [30, 31]), provide a reliable model of social relations in the society. While the previous studies of social networks describe the social relations and links only in the frame work of linear matrix algebra we introduce a nonlinear interaction in such social networks. Similarly to the NLIRM results [17] our studies show that above a certain chaos border for the strength of the nonlinearity dynamical thermalization takes place in social networks being well described by the steady-state Rayleigh-Jeans distribution. This gives an additional support to the Wealth Thermalization Hypothesis. In this way the problem of emergence of statistical laws from dynamical equations of motion finds new application perspectives.
The main results of this research are presented in Part I for links between wealth inequality and Rayleigh-Jeans thermalization and in Part II for dynamical thermalization in social networks supporting the Wealth Thermalization Hypothesis discussed in Part I. Figures in there are marked as Fig.IXX and Fig.IIXX respectively. Additional material and Figures are presented in Appendix A related to Part I and Appendix B related to Part II, in the Appendix A and B Figures are marked as Fig. AXX and Fig.BXX omitting the word Appendix A or B.
Part I: Wealth Thermalization Hypothesis
I.1 Prologue I
@@
The wealth distribution in the human society is characterized by a striking inequality (see e.g. [32, 33, 34]). Thus for the whole world 50% of the population owns only 2% of total wealth, while 10% of population owns 75% of total wealth and 1% of population owns 38% of total wealth [33].
@@
The distribution of wealth is usually described by the Lorenz curve [35, 34] which gives the dependence of accumulated normalized wealth $0\leq w\leq 1$ on the cumulated normalized fraction of population or households $0\leq h\leq 1$. Thus the equipartition of wealth corresponds to the diagonal $w=h$ and the doubled area between diagonal and the Lorenz curve $w(h)$ determines the Gini coefficient $0\leq G\leq 1$ [36, 34]. Values of $G$ can be found in [37] for world countries in 2021 being in the range $0.59<G<0.90$; for the whole world $G=0.889$.
@@
Diverse methods of statistical mechanics and physical kinetics [24, 23, 38] have been proposed and used by different research groups [39, 40, 41, 42, 43, 44, 45, 46, 47]. Various models of interacting agents are investigated including Random Asset Exchange models [39, 40, 41, 42, 43, 44, 45, 46, 47]. In several of these models there is appearance of some kind of oligarchic phase with a significant wealth accumulation by a group of agents [42, 45, 46, 47]. The specific arguments are presented in a favor of the Boltzmann-Gibbs type description of distribution of money, wealth and income [41, 43]. Also a nonlinear Fokker–Planck description of asset exchange is proposed [45, 46] with emergence of oligarchic phase. A few important elements are stressed in [45, 46]: the conservation of two integrals of system evolution being the total wealth and total norm (or number of agents), the argument in favor of consideration of wealth instead of money based on the small-transaction approximation. The conservation of two integrals is rather natural assumption since a Gross domestic product and population of a country or the whole world are only weakly changed on a typical time scale of one year.
The above models give interesting insights for understanding of certain features of wealth distribution in the world countries but they remain model specific and their universality remains questionable. The universality of the Boltzmann-Gibbs thermal distribution is the ground element of the approach developed in [41, 43] but it does not capture emergence of a huge condensate of poverty in various countries.
@@
Here we assume that the system wealth has certain states $0\leq m<N$ with energies $E_{m}$ and the population probabilities in these states are $\rho_{m}$. Thus a systen wealth is associated with a system energy. Also in (I.1) the parameters $T$ and $\mu(T)$ are the system temperature and its chemical potential dependent on $T$. As in [45] there are two conserved integrals of motion being the total norm of population, fixed to be unity for convenience, $\sum_{m}\rho_{m}=1$, and the system average wealth being its total energy $\sum_{m}E_{m}\rho_{m}=E$. For a given system energy $E$ and unity norm these two integrals of motion determine the system temperature $T$ and its chemical potential $\mu(T)$. The entropy $S$ of the system is determined by the usual relation [24, 23]: $S=-\sum_{m}\rho_{m}\ln\rho_{m}$ with the implicit theoretical dependencies on temperature $E(T)$, $S(T)$, $\mu(T)$.